is it possible that 4×10-19 cb charge is present on any body?
 

Grade:12

2 Answers

Arun
4 years ago
Dear student
 
As we know
 
Q = ne
 
4* 10^-19 = n * 1.6 * 10^-19
 
n = 4/1.6
 
As this is not an integer. Hence this cam not be possible
Vikas TU
4 years ago